# SeedVR2: One-Step Video Restoration via Diffusion Adversarial Post-Training

### v1

@@ Abstract
Work was done during Jianyi Wang’s internship at ByteDance Seed (iceclearwjy@gmail.com).
Recent advances in diffusion-based video restoration (VR) demonstrate significant improvement in visual quality, yet yield a prohibitive computational cost during inference. While several distillation-based approaches have exhibited the potential of one-step image restoration, extending existing approaches to VR remains challenging and underexplored, particularly when dealing with high-resolution video in real-world settings. In this work, we propose a one-step diffusion-based VR model, termed as SeedVR2, which performs adversarial VR training against real data. To handle the challenging high-resolution VR within a single step, we introduce several enhancements to both model architecture and training procedures. Specifically, an adaptive window attention mechanism is proposed, where the window size is dynamically adjusted to fit the output resolutions, avoiding window inconsistency observed under high-resolution VR using window attention with a predefined window size. To stabilize and improve the adversarial post-training towards VR, we further verify the effectiveness of a series of losses, including a proposed feature matching loss without significantly sacrificing training efficiency. Extensive experiments show that SeedVR2 can achieve comparable or even better performance compared with existing VR approaches in a single step.
@@ 1 Introduction
figure Speed and performance comparisons. Our SeedVR2 demonstrates impressive restoration capabilities, offering fine details and enhanced visual realism. While achieving comparable performance with SeedVR [67], our SeedVR2 is over 4×4$\times$4 × faster than existing diffusion-based video restoration approaches [97, 79, 14, 74] (We use 50 sampling steps for these baselines to maintain stable performance), even with at least four times the parameter count (Zoom-in for best view).
@@
Diffusion models [38, 16, 51, 62] are becoming the the de-facto model for real-world image restoration (IR) [68, 84, 36, 87, 85, 88, 96] and video restoration (VR) [97, 74, 79, 67, 30]. Though these approaches show promise in generating realistic details, they typically require tens of steps to generate a video sample, leading to considerably high computational cost and latency. Such significant cost is further amplified when processing long videos at high resolutions.
@@
Inspired by recent advances in diffusion acceleration [56, 82, 44, 42], several one-step diffusion IR approaches [71, 54, 99, 27, 10, 48, 75, 15, 28, 90, 85, 72] have been proposed, showing potential in generating promising results comparable to that of multi-step approaches. The majority of these methods [71, 54, 99, 27, 10, 48, 75, 15, 72] rely on distillation from a pre-trained teacher model, suffering from an undesired upper bound constrained by the teacher model. The high computational cost of the teacher model further makes it less practical to apply these methods to VR. The closest to our work are recent distillation-free one-step IR methods that either learn from a discriminator prior [28] or a generative prior [90, 85]. These methods save computational cost by training on an implicit teacher model, i.e., diffusion prior [55, 49] with LoRA [18]. Given the limited capability of existing video diffusion as prior, our work turns to explore one-step VR without depending on any teacher models or frozen prior, avoiding introducing the possible bias learned by these models.
@@
Achieving one-step VR, especially under high resolutions, is challenging, yet underexplored. In this paper, we introduce a new method, SeedVR2, for one-step VR towards real-world scenarios. Our method follows Adversarial Post-Training (APT) [34] to adopt a pre-trained diffusion transformer, i.e., SeedVR [67] as initialization, and continues to fully tune the whole network using the adversarial training objective against real data. Compared with previous one-step IR methods, SeedVR2 eliminates the substantial cost associated with pre-computing video samples from the diffusion teacher during distillation. Moreover, without the constraint from a diffusion teacher or prior, SeedVR2 presents the potential to surpass the initial model, demonstrating comparable or even superior performance to multi-step VR diffusion models.
@@
While it is applicable to directly adopt APT [34] for VR, we empirically observe several key aspects that can be improved based on the nature of VR. First, given the low-quality input as a condition, we observe a more stable training process of VR compared with text-to-video generation [34], i.e., no obvious mode collapse is observed with only a single stage of adversarial training. However, we notice a performance drop when handling heavy degradations. We hereby adopt a progressive distillation [53] before the adversarial training to maintain the restoration capability under one-step generation. Second, when applying window attention with a predefined window size on high-resolution VR, e.g., over 2K resolution, we observe visible boundary artifacts between window patches. We conjecture this is due to the improper settings of the window size and training resolutions, e.g., too large window sizes compared with relatively small training resolutions, making the model insufficiently trained on handling window shifting. Such a predefined window manner may further limit the robustness of 3D Rotary Positional Embedding (RoPE) [63] inside each window when dealing with inputs with various resolutions. To tackle this problem, we propose an adaptive window attention mechanism to dynamically adjust the window size within a certain range, significantly improving the robustness of the model when handling arbitrary-resolution inputs. Third, adversarial training with the exceptionally large generator and discriminator can be unstable even with APT, i.e., a performance deterioration can be observed after long training, e.g., 20k iterations. We follow Huang et al. [19] to enhance the training stability by introducing RpGAN [20] and an additional approximated R2 regularization loss. While L1 loss and LPIPS loss [92] are commonly used in VR training for better perception-distortion tradeoff [2], the necessity to calculate LPIPS in pixel space makes it unaffordable for high-resolution video training. Training a latent LPIPS model [21] is also not applicable due to the lack of video-specific data. We instead propose a feature matching loss to replace the LPIPS loss for efficient adversarial training. Specifically, we directly extract multiple features from different layers of the discriminator and measure the feature distance between the prediction and ground-truth. We empirically show that such a feature matching loss is an effective alternative in our case.
@@
To our knowledge, SeedVR2 is among the first to demonstrate the feasibility of one-step video restoration or super-resolution using a diffusion transformer. Benefiting from the adversarial training with specific designs for VR, we are able to train the largest-ever VR GAN ($\sim$16B for the generator and discriminator in total), which can achieve high-quality restoration in a single sampling step with high efficiency.
The main contributions of our work are as follows:
@@ 2 Related Work
Video Restoration. Traditional video restoration (VR) methods  [4, 5, 31, 32, 26, 8, 83, 69] primarily concentrate on synthetic datasets, suffering from limited effectiveness in real-world scenarios. More recent efforts [6, 73, 95] have shifted focus towards real-world scenarios, but still struggle with generating realistic textures due to constrained generative capabilities. Inspired by the rapid progress in diffusion models [59, 16, 78, 47, 51, 57], several diffusion-based VR methods [97, 14, 79, 74, 30] have emerged, demonstrating remarkable performance. While fine-tuning on a diffusion prior [51, 94] improves efficiency, these methods still inherit the inherent limitations of the diffusion prior, i.e., inefficient autoencoder and inflexible resolution scalability as discussed by Wang et al. [67]. The most recent work [67] proposes to fully train a large diffusion transformer model with a shifted window attention and a casual video autoencoder, achieving impressive performance with relatively high efficiency. However, the need for tens of steps to sample a video still leads to unfriendly latency in real-world applications. By introducing APT [34] into diffusion-based VR, our approach is capable of achieving one-step VR with high quality, which, to the best of our knowledge, is among the earliest explorations of one-step diffusion-based VR.
@@
Diffusion Acceleration. As discussed by Lin et al. [34], most of the existing approaches either distill the deterministic probability flow learned by a diffusion teacher model using fewer steps (i.e., deterministic methods) or approximate the same distribution of a diffusion teacher model (i.e., distributional methods). Specifically, deterministic methods include progressive distillation [53], consistency distillation [61, 60, 40, 41, 42], and rectified flow [37, 39, 77]. Though these methods can be easily trained with simple regression loss, blurry results can be commonly observed with very few steps, i.e., less than 8 steps [61, 41, 42]. In addition to directly predicting the outputs of the teacher model, distributional methods turn to adversarial training [55, 76, 7, 22, 44], score distillation [82, 43], both [81, 56, 3], and combining with deterministic methods [24, 33, 50] to resemble the distribution of a teacher model. Most recent approaches [76, 34] instead directly fine-tune a pre-trained diffusion model on real data with adversarial training, leading to superior performance with one-step sampling. While several acceleration approaches [35, 65, 89, 34] have been extended to video generation, the one-step acceleration for video diffusion restoration is still underexplored, inspiring us to make an early attempt in this direction.
@@
One-step Restoration. While conventional GAN-based real-world restoration approaches [91, 70, 6, 95] can achieve one-step restoration, their poor generation ability usually leads to suboptimal results. To improve the sampling efficiency of diffusion-based approaches [68, 84, 97, 79], ResShift [87, 88] shifts the initial sampling distribution from a standard Gaussian distribution to the distribution of low-quality images, achieving a fast sampling of up to 4 steps. Recent advances further achieve one-step sampling via distillation [71, 54, 99, 27, 10, 48, 75, 15], adversarial training [28], or tuning on a prior with additional trainable layers [90, 85, 72]. However, all these methods focus on image restoration and may not be suitable for VR due to the lack of temporal design and unsatisfactory generation quality. Compared with these methods, our method achieves one-step VR with substantially better quality, especially under high-resolution real-world scenarios.
@@ 3 Methodology
The objective of SeedVR2 is to perform one-step Video Restoration (VR) by upscaling an input video into a high-resolution output. SeedVR2 builds upon previous works [67, 34], with preliminary concepts introduced in Sec.3.1.
@@ 3.1 Preliminaries: Diffusion Adversarial Post-Training
Diffusion Adversarial Post-Training (APT) [34] is a diffusion acceleration approach that converts a multi-step diffusion model to a one-step generator. There are mainly two training stages in APT, i.e., deterministic distillation and Adversarial Post-Training (APT). During the deterministic distillation, a distilled model is first trained following discrete-time consistency distillation [61, 60] with mean squared error loss. The teacher model generates distillation supervision with a constant classifier-free guidance [17] scale of 7.5 and a predefined negative prompt. As for adversarial training, the discriminator is first initialized by the pre-trained diffusion network, and then additional cross-attention-only transformer blocks are introduced to generate logits for loss calculation. To stabilize the adversarial training while avoiding higher-order gradient computation, APT proposes an approximated R1 loss [52] to regularize the discriminator, and the final loss for the discriminator is a non-saturating GAN loss [11] combined with the approximated R1 loss. Our method employs a similar network architecture to APT, where both the generator and discriminator are diffusion transformers, as shown in Figure 3.2.
@@ 3.2 Adaptive Window Attention
To improve the robustness of window attention for high-resolution inputs with arbitrary sizes, we propose an adaptive window attention mechanism that allows the window size to be dynamically adjusted to fit the input resolution, as shown in Figure 3.2. During training, given a video feature $X\in\mathbb{R}^{d_{t}\times d_{h}\times d_{w}\times d_{c}}$, where $d_{h}\times d_{w}=45\times 80$ (i.e., the feature resolution under 720p), the window size of our attention is calculated accordingly as follows:
@@
[…]
figure Model architecture and the partition of the adaptive attention window. We improve the Swin-MMDIT [67] with an adaptive window partition, i.e., the window size is ensured via a $3\times 3$ partition on the resized LQ input ($\rm{Height}\times\rm{Width}=960\times 960$). The features for calculating the feature matching loss are extracted before the cross-attention layers used in APT [34].
@@
where $d_{h}\times d_{w}=45\times 80$ is the training resolution. This ensures $\frac{\tilde{d}_{h}}{\tilde{d}_{w}}=\frac{\hat{d}_{h}}{\hat{d}_{w}}$ and $\tilde{d}_{h}\times\tilde{d}_{w}=d_{h}\times d_{w}$. The final window size for test-time attention is then obtained by substituting $(d_{t},{\color[rgb]{0,0,1}d_{h}},{\color[rgb]{0,0,1}d_{w}})$ in Eq. (1) with $(\hat{d}_{t},{\color[rgb]{0,0,1}\tilde{d}_{h}},{\color[rgb]{0,0,1}\tilde{d}_{w}})$. This adaptive partition strategy enhances consistency between training and testing configurations and substantially alleviates boundary artifacts in high-resolution predictions, as illustrated in Figure 4.2.

### v2

@@ Abstract
Recent advances in diffusion-based video restoration (VR) demonstrate significant improvement in visual quality, yet yield a prohibitive computational cost during inference. While several distillation-based approaches have exhibited the potential of one-step image restoration, extending existing approaches to VR remains challenging and underexplored, particularly when dealing with high-resolution video in real-world settings. In this work, we propose a one-step diffusion-based VR model, termed as SeedVR2, which performs adversarial VR training against real data. To handle the challenging high-resolution VR within a single step, we introduce several enhancements to both model architecture and training procedures. Specifically, an adaptive window attention mechanism is proposed, where the window size is dynamically adjusted to fit the output resolutions, avoiding window inconsistency observed under high-resolution VR using window attention with a predefined window size. To stabilize and improve the adversarial post-training towards VR, we further verify the effectiveness of a series of losses, including a proposed feature matching loss without significantly sacrificing training efficiency. Extensive experiments show that SeedVR2 can achieve comparable or even better performance compared with existing VR approaches in a single step. Work was done during Jianyi Wang’s internship at ByteDance Seed (iceclearwjy@gmail.com). Now at Apple.
@@ 1 Introduction
Figure 1: Speed and performance comparisons. Our SeedVR2 demonstrates impressive restoration capabilities, offering fine details and enhanced visual realism. While achieving comparable performance with SeedVR (Wang et al., 2025), our SeedVR2 is over $4\times$ faster than existing diffusion-based video restoration approaches (Zhou et al., 2024, Yang et al., 2024, He et al., 2024a, Xie et al., 2025) (We use 50 sampling steps for these baselines to maintain stable performance), even with at least four times the parameter count (Zoom-in for best view).
@@
Diffusion models (Liu et al., 2023, Ho et al., 2020, Rombach et al., 2022, Song et al., 2021) are becoming the the de-facto model for real-world image restoration (IR) (Wang et al., 2024b, Yu et al., 2024, Lin et al., 2024b, Yue et al., 2023; 2025; 2024) and video restoration (VR) (Zhou et al., 2024, Xie et al., 2025, Yang et al., 2024, Wang et al., 2025, Li et al., 2025b). Though these approaches show promise in generating realistic details, they typically require tens of steps to generate a video sample, leading to considerably high computational cost and latency. Such significant cost is further amplified when processing long videos at high resolutions.
@@
Inspired by recent advances in diffusion acceleration (Sauer et al., 2024b, Yin et al., 2024b, Luo et al., 2025; 2023b), several one-step diffusion IR approaches (Wang et al., 2024c, Zhu et al., 2024, Li et al., 2025a, Dong et al., 2025, Xie et al., 2024, Li et al., 2024, Yue et al., 2025, Wu et al., 2024) have been proposed, showing potential in generating promising results comparable to that of multi-step approaches. The majority of these methods (Wang et al., 2024c, Zhu et al., 2024, Li et al., 2025a, Dong et al., 2025, Xie et al., 2024, Wu et al., 2024) rely on distillation from a pre-trained teacher model, suffering from an undesired upper bound constrained by the teacher model. The high computational cost of the teacher model further makes it less practical to apply these methods to VR. The closest to our work are recent distillation-free one-step IR methods that either learn from a discriminator prior (Li et al., 2024) or a generative prior (Zhang et al., 2024, Yue et al., 2025). These methods save computational cost by training on an implicit teacher model, i.e., diffusion prior (Sauer et al., 2024a, Podell et al., 2024) with LoRA (Hu et al., 2022a). Given the limited capability of existing video diffusion as prior, our work turns to explore one-step video restoration via adversarial training without frozen priors, making it possible to alleviate the bias learned by these models.
@@
Achieving one-step VR, especially under high resolutions, is challenging, yet underexplored. In this paper, we introduce a new method, SeedVR2, for one-step VR towards real-world scenarios. Our method follows Adversarial Post-Training (APT) (Lin et al., 2025) to adopt a pre-trained diffusion transformer, i.e., SeedVR (Wang et al., 2025) as initialization, and continues to fully tune the whole network using the adversarial training objective against real data. Compared with previous one-step IR methods, SeedVR2 eliminates the substantial cost associated with pre-computing video samples from the diffusion teacher during distillation. Moreover, without the constraint from a diffusion teacher or prior, SeedVR2 presents the potential to surpass the initial model, demonstrating comparable or even superior performance to multi-step VR diffusion models.
@@
While it is applicable to directly adopt APT for VR, we empirically observe several key aspects that can be improved based on the nature of VR. First, given the low-quality input as a condition, we observe a more stable training process of VR compared with text-to-video generation (Lin et al., 2025), i.e., no obvious mode collapse is observed with only a single stage of adversarial training. However, we notice a performance drop when handling heavy degradations. We hereby adopt a progressive distillation (Salimans and Ho, 2022) before the adversarial training to maintain the restoration capability under one-step generation. Second, when applying window attention with a predefined window size on high-resolution VR, e.g., over 2K resolution, we observe visible boundary artifacts between window patches. We conjecture this is due to the improper settings of the window size and training resolutions, e.g., too large window sizes compared with relatively small training resolutions, making the model insufficiently trained on handling window shifting. Such a predefined window manner may further limit the robustness of 3D Rotary Positional Embedding (RoPE) (Su et al., 2024) inside each window when dealing with inputs with various resolutions. To tackle this problem, we propose an adaptive window attention mechanism to dynamically adjust the window size within a certain range, significantly improving the robustness of the model when handling arbitrary-resolution inputs. Third, adversarial training with the exceptionally large generator and discriminator can be unstable even with APT, i.e., a performance deterioration can be observed after long training, e.g., 20k iterations. We follow  Huang et al. (2024) to enhance the training stability by introducing RpGAN (Jolicoeur-Martineau, 2019) and an additional approximated R2 regularization loss. While L1 loss and LPIPS loss (Zhang et al., 2018a) are commonly used in VR training for better perception-distortion tradeoff (Blau and Michaeli, 2018), the necessity to calculate LPIPS in pixel space makes it unaffordable for high-resolution video training. Training a latent LPIPS model (Kang et al., 2024a) is also not applicable due to the lack of video-specific data. We instead propose a feature matching loss to replace the LPIPS loss for efficient adversarial training. Specifically, we directly extract multiple features from different layers of the discriminator and measure the feature distance between the prediction and ground-truth. We empirically show that such a feature matching loss is an effective alternative in our case.
@@
To our knowledge, SeedVR2 is among the early attempts to demonstrate the feasibility of one-step video restoration or super-resolution using a diffusion transformer. Benefiting from the adversarial training with specific designs for VR, we are able to train the largest-ever VR GAN ($\sim$16B for the generator and discriminator in total), which can achieve high-quality restoration in a single sampling step with high efficiency. The main contributions of our work are as follows:
@@ 2 Related Work
Video Restoration. Traditional video restoration (VR) methods  (Chan et al., 2021; 2022b, Liang et al., 2024; 2022, Li et al., 2023, Chen et al., 2024, Youk et al., 2024, Wang et al., 2019) primarily concentrate on synthetic datasets, suffering from limited effectiveness in real-world scenarios. More recent efforts (Chan et al., 2022a, Xie et al., 2023, Zhang and Yao, 2024) have shifted focus towards real-world scenarios, but still struggle with generating realistic textures due to constrained generative capabilities. Inspired by the rapid progress in diffusion models (Sohl-Dickstein et al., 2015, Ho et al., 2020, Yang et al., 2021, Nichol et al., 2022, Rombach et al., 2022, Seawead et al., 2025), several diffusion-based VR methods (Zhou et al., 2024, He et al., 2024a, Yang et al., 2024, Xie et al., 2025, Li et al., 2025b) have emerged, demonstrating remarkable performance. While fine-tuning on a diffusion prior (Rombach et al., 2022, Zhang et al., 2023) improves efficiency, these methods still inherit the inherent limitations of the diffusion prior, i.e., inefficient autoencoder and inflexible resolution scalability as discussed by Wang et al. (2025). The most recent work (Wang et al., 2025) proposes to fully train a large diffusion transformer model with a shifted window attention and a casual video autoencoder, achieving impressive performance with relatively high efficiency. However, the need for tens of steps to sample a video still leads to unfriendly latency in real-world applications. By introducing APT (Lin et al., 2025) into diffusion-based VR, our approach is capable of achieving one-step VR with high quality, which, to the best of our knowledge, is among the earliest explorations of one-step diffusion-based VR.
@@
Diffusion Acceleration. As discussed by Lin et al. (2025), most of the existing approaches either distill the deterministic probability flow learned by a diffusion teacher model using fewer steps (i.e., deterministic methods) or approximate the same distribution of a diffusion teacher model (i.e., distributional methods). Specifically, deterministic methods include progressive distillation (Salimans and Ho, 2022), consistency distillation (Song et al., 2023, Song and Dhariwal, 2024, Lu and Song, 2025, Luo et al., 2023a; b), and rectified flow (Liu et al., 2023; 2024, Yan et al., 2024). Though these methods can be easily trained with simple regression loss, blurry results can be commonly observed with very few steps, i.e., less than 8 steps (Song et al., 2023, Luo et al., 2023a; b). In addition to directly predicting the outputs of the teacher model, distributional methods turn to adversarial training (Sauer et al., 2024a, Xu et al., 2024, Chen et al., 2025a, Kang et al., 2024b, Luo et al., 2025), score distillation (Yin et al., 2024b, Luo et al., 2023c), both (Yin et al., 2024a, Sauer et al., 2024b, Chadebec et al., 2025), and combining with deterministic methods (Kohler et al., 2024, Lin et al., 2024a, Ren et al., 2024) to resemble the distribution of a teacher model. Most recent approaches (Xu et al., 2024, Lin et al., 2025) instead directly fine-tune a pre-trained diffusion model on real data with adversarial training, leading to superior performance with one-step sampling. While several acceleration approaches (Lin and Yang, 2024, Wang et al., 2024a, Zhai et al., 2024, Lin et al., 2025) have been extended to video generation, the one-step acceleration for video diffusion restoration is still underexplored, inspiring us to make an early attempt in this direction.
@@
One-step Restoration. While conventional GAN-based real-world restoration approaches (Zhang et al., 2021, Wang et al., 2021, Chan et al., 2022a, Zhang and Yao, 2024, Zhou et al., 2022) can achieve one-step restoration, their poor generation ability usually leads to suboptimal results. To improve the sampling efficiency of diffusion-based approaches (Wang et al., 2024b, Yu et al., 2024, Zhou et al., 2024, Yang et al., 2024), ResShift (Yue et al., 2023; 2024) shifts the initial sampling distribution from a standard Gaussian distribution to the distribution of low-quality images, achieving a fast sampling of up to 4 steps. Recent advances further achieve one-step sampling via distillation (Wang et al., 2024c, Sami et al., 2024, Zhu et al., 2024, Li et al., 2025a, Dong et al., 2025, Noroozi et al., 2024, Xie et al., 2024, He et al., 2024b), adversarial training (Li et al., 2024), or tuning on a prior with additional trainable layers (Zhang et al., 2024, Yue et al., 2025, Wu et al., 2024). However, all these methods focus on image restoration and may not be suitable for VR due to the lack of temporal design and unsatisfactory generation quality. Compared with these methods, our method achieves one-step VR with substantially better quality, especially under high-resolution real-world scenarios.
@@ 3 Methodology
The objective of SeedVR2 is to perform one-step Video Restoration (VR) by upscaling an input video into a high-resolution output. SeedVR2 builds upon previous works (Wang et al., 2025, Lin et al., 2025), with preliminary concepts introduced in Sec.3.1.
@@ 3.1 Preliminaries: Diffusion Adversarial Post-Training
Diffusion Adversarial Post-Training (APT) (Lin et al., 2025) is a diffusion acceleration approach that converts a multi-step diffusion model to a one-step generator. There are mainly two training stages in APT, i.e., deterministic distillation and Adversarial APT. During the deterministic distillation, a distilled model is first trained following discrete-time consistency distillation (Song et al., 2023, Song and Dhariwal, 2024) with mean squared error loss. The teacher model generates distillation supervision with a constant classifier-free guidance (Ho and Salimans, 2021) scale of 7.5 and a predefined negative prompt. As for adversarial training, the discriminator is first initialized by the pre-trained diffusion network, and then additional cross-attention-only transformer blocks are introduced to generate logits for loss calculation. To stabilize the adversarial training while avoiding higher-order gradient computation, APT proposes an approximated R1 loss (Roth et al., 2017) to regularize the discriminator, and the final loss for the discriminator is a non-saturating GAN loss (Goodfellow et al., 2014) combined with the approximated R1 loss. Our method employs a similar network architecture to APT, where both the generator and discriminator are diffusion transformers, as shown in Figure 2.
@@ 3.2 Adaptive Window Attention
To improve the robustness of window attention for high-resolution inputs with arbitrary sizes, we propose an adaptive window attention mechanism that allows the window size to be dynamically adjusted to fit the input resolution, as shown in Figure 2. During training, given a video feature $X\in\mathbb{R}^{d_{t}\times d_{h}\times d_{w}\times d_{c}}$, where $d_{h}\times d_{w}=45\times 80$ (i.e., the feature resolution under 720p), the window size of our attention is calculated accordingly as follows:
@@
where $n_{t}$, $n_{h}$ and $n_{w}$ decide the number of windows along dimension $d_{t}$, $d_{h}$ and $d_{w}$, respectively. The ceiling function is represented as $\left\lceil\cdot\right\rceil$, and the term $\min(d_{t},30)$ sets an upper bound to $d_{t}$ to avoid the gap of sequence length between training and inference. Note that although the resolutions of our training data are around 720p, the aspect ratio of width and height can vary a lot, leading to various window sizes during training. Such a design ensures a better generalization ability toward inputs of different resolutions with diverse window sizes.
Figure 2: Model architecture and the adaptive attention window. We improve the Swin-MMDIT (Wang et al., 2025) with an adaptive window partition, i.e., the window size is ensured via a $3\times 3$ partition on the resized LQ input ($\rm{Height}\times\rm{Width}=960\times 960$). The features for calculating the feature matching loss are extracted before the cross-attention layers used in APT (Lin et al., 2025).
@@
where $d_{h}\times d_{w}=45\times 80$ is the training resolution. This ensures $\frac{\tilde{d}_{h}}{\tilde{d}_{w}}=\frac{\hat{d}_{h}}{\hat{d}_{w}}$ and $\tilde{d}_{h}\times\tilde{d}_{w}=d_{h}\times d_{w}$. The final window size for test-time attention is then obtained by substituting $(d_{t},{\color[rgb]{0,0,1}d_{h}},{\color[rgb]{0,0,1}d_{w}})$ in Eq. (1) with $(\hat{d}_{t},{\color[rgb]{0,0,1}\tilde{d}_{h}},{\color[rgb]{0,0,1}\tilde{d}_{w}})$. This adaptive partition strategy enhances consistency between training and testing configurations and substantially alleviates boundary artifacts in high-resolution predictions, as illustrated in Figure 4.